### v1

# Are ECGs enough? Deep learning classification of cardiac anomalies using only electrocardiograms

Abstract
Electrocardiography (ECG) is an essential tool for diagnosing multiple cardiac anomalies: it provides valuable clinical insights, while being affordable, fast and available in many settings. However, in the current literature, the role of ECG analysis is often unclear: many approaches either rely on additional imaging modalities, such as Computed Tomography Pulmonary Angiography (CTPA), which may not always be available, or do not effectively generalize across different classification problems. Furthermore, the availability of public ECG datasets is limited and, in practice, these datasets tend to be small, making it essential to optimize learning strategies. In this study, we investigate the performance of multiple neural network architectures in order to assess the impact of various approaches. Moreover, we check whether these practices enhance model generalization when transfer learning is used to translate information learned in larger ECG datasets, such as PTB-XL and CPSC18, to a smaller, more challenging dataset for pulmonary embolism (PE) detection. By leveraging transfer learning, we analyze the extent to which we can improve learning efficiency and predictive performance on limited data.
Keywords:
Deep Learning Electrocardiography Anomaly detection.
1 Introduction
Cardiovascular diseases (CVDs) represent a significant global health challenge, being a leading cause of mortality and a major contributor to reduced quality of life [12]. Electrocardiography (ECG) is a fundamental tool in diagnosing and managing various cardiac and non-cardiac conditions that is widely utilized in both pre-hospital and in-hospital settings due to its affordability, speed, and accessibility [21].
Despite technological advancements that have simplified ECG-based diagnostics, there is still substantial room for improvement [16]. For instance, PE is one of the deadliest cardiovascular emergencies, yet ECGs alone are insufficient for reliable detection due to their non-specific signs and symptoms. As a result, Computed Tomography Pulmonary Angiography (CTPA), which is not available in every scenario, is often required for definitive diagnosis [8].
The recent rise of foundational AI models has revolutionized multiple fields [14, 18], yet anomaly detection using ECGs in clinical practice remains highly specialized, making generalization difficult. Deep learning has the potential to bridge this gap, but its effectiveness is limited since there are few publicly available ECG datasets. Even worse is that many of these are small (fewer than 10,000 samples) and highly imbalanced, posing significant challenges for model training and generalization.
Given these constraints, our work aims to discover what is the baseline performance of deep learning applied to ECGs and what impact can transfer learning have in enhancing anomaly detection using small, imbalanced datasets. Furthermore, we demonstrate that leveraging this knowledge enhances PE detection using only ECG data and that, despite limited amount of data, clinical decision-making can still be significantly improved.
All code, architectures and hyperparameters are publicly available at https:github.com/joaodsmarques/Are-ECGs-enough-Deep-Learning-Classifiers.
2 Background
The application of deep learning to ECG analysis is not new. Multiple studies have explored this area over the years [6, 24]. However, progress in the field remains challenging, largely due to the limited availability of publicly accessible datasets.
A 1D-CNN is often the standard approach for ECG classification, as convolutional layers effectively capture local features from the signal, achieving high performance in several classification tasks [28]. However, these architectures struggle to recognize long-range dependencies, which are critical for analyzing time-series data such as ECGs. Recurrent neural networks (RNNs), particularly LSTMs and GRUs [4], are well-suited for capturing these temporal dependencies. Consequently, they have also been widely used for ECG classification. As demonstrated by Zahra Ebrahimi et al. [7], both 1D-CNNs and RNN-based models can extract relevant features for ECG classification.
Recently, Hemaxi Narotamo et al. [13] explored multiple deep learning architectures for detecting cardiac anomalies using the PTB-XL dataset. In their 1D experiments, they found that CRNNs, using GRU and LSTM, offer an effective balance between feature extraction and capturing temporal dependencies.
Beyond traditional architectures, new transformer-inspired models have been proposed for ECG analysis. For example, ECG-Mamba [15] achieved an F1-score of 75% on PTB-XL and 79% on CPSC18, demonstrating the potential of new architectures in ECG classification.
Additionally, large-scale foundation models are emerging in ECG research. HuBERT-ECG [5], based on the BERT architecture, is a self-supervised model pre-trained on 9.1 million 12-lead ECGs spanning 164 cardiovascular conditions, showcasing the feasibility of large-scale ECG models.
In the context of pulmonary embolism (PE) classification, Somani et al. demonstrated that ECGs provide valuable diagnostic information for PE detection and can enhance the performance beyond what is achievable using CTPA alone [23].
In this work, we explore multiple deep learning architectures applied to different ECG datasets to assess their effectiveness in learning from raw ECG signals alone, particularly when dealing with limited training data.
3 Methodology
Publicly available datasets for ECG anomaly detection are limited and those that exist are often small and imbalanced. Therefore, it is important to maximize the utility of these datasets to extract meaningful insights that can be leveraged in real-world applications. To achieve this, we guide our experimental evaluation by the following key questions:
  • What performance can we expect when training from scratch? We optimize multiple models for 3 different datasets, in order to assess the raw power of deep learning on those settings
  • How powerful is transfer learning? We explore whether transfer learning from slightly bigger datasets to a smaller setting for PE detection can improve performance.
  • Can our approaches surpass clinical performance in the PE setting? We analyze if our best achieved model can improve medical diagnostic on the PE detection task, when compared to clinical procedures.
3.1 Datasets
We focus this research on 3 datasets: PTB-XL [27], China Physiological Signal Challenge 2018 (CPSC18) [11] and PE-HSM [25].
The selection of datasets is based on the objective of training on the largest and most complete publicly available dataset, PTB-XL, while using CPSC18 as a comparative benchmark due to its relatively large size. Additionally, we evaluate whether the insights gained from these datasets generalize to a smaller, more challenging dataset, PE-HSM, focused on PE detection. Notice that all selected datasets share characteristics, including a sampling frequency of 500 Hz, 12-lead ECG recordings, and adaptable signal lengths, ensuring consistency across experiments. Table 1 provides an overview of these datasets.
Table 1: Overview of the datasets used for each selected task.
Dataset Availability Task Classification Samples Length
PTB-XL Public Diagnostic Multi-label (5) 21,837 10s
CPSC18 Public Arrhythmia Multi-label (9) 6,877 6–60s
PE-HSM Private Acute PE Binary (1) 927 10s
The composition and distribution of each dataset can be found in the publications they were introduced [11, 25, 27]. In the PTB-XL dataset, we focused on the five diagnostic superclasses rather than the 44 diagnostic subclasses. Fine-tuning models for all these subclasses is very challenging due to class imbalance, which can negatively impact model representations, affecting the performance of transfer learning, as the learned representations might be biased toward dominant classes. Figure 1 illustrates the class distributions for the experimental settings.
Figure 1: PTB-XL class distribution (Left) and CPSC18 class distribution (Right).
For PTB-XL, we followed the authors’ recommendations [27], using fold 9 for validation and fold 10 for testing. For CPSC18, we generated 10 folds, ensuring that each fold contains samples from all classes, and performed cross-validation.
Finally, for PE-HSM, we requested access to the dataset from Hospital de Santa Maria in Lisbon and adopted the same training and test splits the authors proposed [25]. The training set consists of 824 samples, including 222 positive samples of PE and 602 negative samples. The test set comprises 103 samples, with 39 positive and 64 negative samples.
3.2 Pre-processing and data augmentation
Pre-processing is crucial for improving model performance, as it helps the models focus on relevant signal features and ignore noise [20]. While specialized algorithms exist for detecting and enhancing specific waveforms, such as R-peaks [19], we adopt a simpler approach to assess the performance of neural networks from raw ECG signals. Our pre-processing pipeline consists of four key steps: noise removal (addressing power line interference and baseline drift), sequence length selection, normalization, and data augmentation. The full pipeline is illustrated in Figure 2.
Figure 2: Pre-processing pipeline.
To mitigate noise, we applied a low-pass filter at 50 Hz to remove power line interference, followed by wavelet decomposition [1]. However, applying this method during training, as required for randomized data augmentation, proved to be computationally expensive. To address this, we instead employed a second-order Butterworth bandpass filter with a high-pass cutoff at 1 Hz for baseline wander suppression and a low-pass cutoff at 45 Hz, which produced similar output quality while being more computationally efficient [13].
For sequence length selection, in order to extract a segment from each ECG lead, we first randomly select a starting point $s$, ensuring that the extracted segment remains within the maximum available length. This process is defined in Equation 1.
$s\sim\mathcal{U}(0,m-l),\quad\text{with}\quad s+l\leq m$ (1)
where $s$ is the randomly chosen starting index, $l$ is the desired segment length, and $m$ is the maximum length of the ECG signal (5000 samples). This ensures that the segment length is consistent across leads and does not exceed the available data range. Sequences exceeding 5000 samples (maximum available length) were truncated, while shorter sequences were zero-padded. A length of 2048 samples was chosen, as it provided similar performance to the longest sequence while being significantly less computationally demanding.
For normalization, we evaluated the impact of various methods, including MinMax, Z-score, Rscale, Logscale, and L2 normalization. Finally, we applied multiple data augmentation techniques, such as flip, random drop, lead drop, square pulse sum and sine sum [17].
3.3 Model architectures
We explored a range of architectures, from simpler models like AlexNet to more complex encoder-based networks, in order to evaluate model robustness, training efficiency and performance.
We started by using AlexNet [9] as a baseline due to its simplicity and fast training speed. Next, we investigated the VGG family [22], identifying VGG11 (with batch normalization) as the most robust contender. We then conducted experiments with ResNet architectures, where ResNet18 emerged as a strong candidate, offering a good balance between simplicity and performance, while also proving to be an effective feature extractor for other deep learning modules.
Beyond CNN-based models, we integrated GRUs and LSTMs [4] to capture temporal dependencies in ECG signals. To further enhance both deep feature extraction and temporal modeling, we employed CRNNs, where ResNet18 with GRU and ResNet18 with LSTM stood out as high-performing combinations.
Given the importance of attention mechanisms in capturing long-range dependencies [26], we studied adding attention to the ResNet18 [25], naming this architecture AttResNet.
Finally, we explored transformer-based models [26], including transformer encoders and Residual Transformers, to assess how larger, more recent architectures perform in this setting.
3.4 Metrics
Given that our task is predominantly multi-label and that class distributions are imbalanced, relying only on accuracy or area under the curve (AUC) can be misleading, as models may achieve high scores by prioritizing the most abundant classes. To provide a more comprehensive evaluation, we additionally report F1-score and macro average precision (MAP), which can be interpreted as an AUC-like metric for the precision-recall curve. Furthermore, in clinical applications, it is critical to minimize false negatives (ensuring high sensitivity) while also keeping high specificity to confirm that detected negatives are truly negative cases. To measure this balance, we also compute the G-mean (GM), defined as:
$G_{\text{mean}}=\sqrt{\text{sensitivity}\times\text{specificity}}$ (2)
4 Results and Discussion
All models were initialized using standard procedures [3] and trained on an NVIDIA V100S GPU (32GB) installed in a DELL PowerEdge C41402 server. To determine the optimal hyperparameter configuration for each architecture, we employed Bayesian Optimization with Optuna [2], running 100 trials with 50 epochs each, with the objective of maximizing the F1-score.
For the loss function, we analyzed both Binary Cross Entropy (BCE) with class weighting, adjusted according to dataset distribution, and Focal Loss [10] with $\gamma=2$ and $\alpha=0.7$. The Focal Loss proved to be the most effective, being considered essential for achieving strong results. The best results obtained on the publicly available datasets, when training from scratch, are summarized in Table 2.
Table 2: Performance (%) of multiple models trained from scratch on PTB-XL and CPSC18 datasets.
PTB-XL CPSC18
Models Acc F1 MAP GM Acc F1 MAP GM
AlexNet 85.9 74.3 76.1 83.8 93.2 74.8 72.4 86.2
VGG11_bn 86.2 74.1 77.3 83.2 94.0 75.0 70.9 86.1
ResNet18 86.8 75.1 78.9 83.9 93.9 75.3 72.0 87.2
CRNNLSTM 87.1 75.8 76.7 84.6 95.1 79.3 76.8 88.4
CRNNGRU 87.2 76.1 76.7 84.9 95.1 79.5 78.3 88.6
AttResNet 87.0 75.6 77.1 85.0 94.8 78.5 76.8 88.1
Transformer 75.0 50.7 44.0 64.9 87.9 57.5 51.5 74.4
ResTransformer 87.2 75.7 76.6 84.0 94.8 78.8 77.5 89.0
The optimized architectures include the following configurations: (a) CNNs: logscale normalization and learning rate (lr) $\approx$ 0.0002; (b) CRNNs: ResNet18, L2 normalization, 2 recurrent layers, 256 hidden size and lr $\approx$ 0.0005; (c) AttResNet: ResNet18 combined with multihead attention layer, 512 embedding dimension, 4 heads, L2 normalization and lr $\approx$ 0.0002; (d) Transformer: 1 convolutional layer to increase dimensions, 4 transformer encoder layers, 512 embedding dimension, 4 heads, zscore normalization and lr $\approx$ 0.0001; (e) ResTransformer: ResNet18, transformer encoder, 512 embedding dimension, 4 heads, zscore normalization and lr $\approx$ 0.0002.
Based on the results obtained in Table 2, we chose 4 models to apply to the PE-HSM dataset: the best CNN (ResNet18), the best CRNN (CRNNGRU), the Attention ResNet (AttResNet) and the ResTransformer. The performance of these 4 models was analyzed both when using transfer learning and training from scratch. Regarding the transfer learning setting, we noted that attention related models (ResTransformer and AttResNet) performed better when all the weights were updated, while the other two performed better when only the classification layer was updated. All the results for this setting are illustrated in Table 3.
Table 3: Performance (%) of the best models on the PE-HSM dataset, both trained from scratch and pretrained.
PE-HSM
Models Pretrain Acc F1 MAP GM
ResNet18 None 58.3 56.6 42.2 59.6
PTB-XL 63.1 61.2 50.8 64.9
CPSC18 63.0 62.4 55.1 62.7
CRNNGRU None 66.0 61.9 48.2 70.1
PTB-XL 68.9 65.4 49.6 72.7
CPSC18 67.0 63.2 54.1 70.81
AttResNet None 64.1 60.2 52.4 67.8
PTB-XL 68.0 64.1 53.8 71.9
CPSC18 64.1 63.4 55.7 63.5
ResTransformer None 66.0 57.6 49.3 72.3
PTB-XL 63.1 60.8 50.3 61.8
CPSC18 68.0 66.7 51.9 68.7
While no model outperforms all others across all metrics, several insights can be obtained from the results. Transfer learning increases performance on small datasets, as evidenced by improvements in both pretrained settings. CRNNs offer the best balance between ease of training and performance, while ResNet18 proves to be a strong and stable feature extractor. Attention modules benefit greatly from feature extractors, whereas transformers struggle due to limited training data.
In the PTB-XL pretraining setting, AttResNet comes close to the CRNN, suggesting that with more data, attention mechanisms might outperform RNNs. Interestingly, ResTransformer performs better when pretrained on CPSC18, than on PTB-XL, likely due to CPSC18’s focus on arrhythmias. Particularly, CPSC18 has a class for Right Bundle Branch Block (RBBB), which is a key ECG indicator for pulmonary embolism: this demonstrates effective domain adaptation for the ResTransformer.
Finally, we evaluated our best model, CRNNGRU (pretrained on PTB-XL) on the PE-HSM test set against two clinical scores for PE diagnosis, Wells score and PEGeD [25], as shown in Figure 3.
Figure 3: CRNNGRU diagnostic performance against PEGeD (Left) and Wells (Right), which are some of the best clinical scores performed by doctors to detect PE [25].
Although both clinical scores are way more sensitive, the CRNNGRU demonstrates superior AUC, specificity and precision, showing that deep learning can be useful in detecting anomalies using only ECGs.
5 Conclusions
In this study, we present a novel methodology for ECG anomaly detection, explaining key steps such as optimal length selection, model-specific normalization, and best practices. We validate this approach across 3 datasets and apply transfer learning to enhance PE diagnosis. Our findings demonstrate that ECGs can be enough, establishing strong classification baselines and contributing to improved clinical decision-making, even in small, imbalanced dataset scenarios.
6 Acknowledgements
This work was supported by project nº 62 - Center for Responsible AI ref. C628696807-00454142, financed by the Recovery and Resilience Facility - Component 5, included in the NextGenerationEU funding program, by project PRELUNA, grant PTDC/CCIINF/4703/2021, and by national funds through FCT, Fundação para a Ciência e a Tecnologia, under the project UIDB/50021/2020 (DOI:10.54499/UIDB/50021/2020). We also extend our thanks to Beatriz Valente Silva, Miguel Nobre Menezes, and Fausto J. Pinto for their research contributions and for granting us access to the PE-HSM dataset.